OpenStudy (anonymous):

Math question, please help. 9|9 – 8x| = 2x + 3 I have to find the extraneous solutions

OpenStudy (anonymous):

not sure what "extraneous solution" means there are two solutions

OpenStudy (anonymous):

if \(x>\frac{9}{8}\) solve \[9(8x-9)=2x+3\]

OpenStudy (anonymous):

Solve the equation. Check for extraneous solutions. 9|9 – 8x| = 2x + 3 (2 points) this is all the question says.

OpenStudy (anonymous):

oh "check" not find

OpenStudy (anonymous):

this means check your answer that is all

OpenStudy (anonymous):

yeah, sorry I didn't realize that until I looked back, but I still don't know how to solve it :/ I am horrible at math.

OpenStudy (anonymous):

solve two equations \[9(9-8x)=2x+3\] and \[9(8x-9)=2x+3\]

OpenStudy (anonymous):

you will get two different solutions, and the "check" part means check your answer to make sure it works can you solve these?

OpenStudy (anonymous):

Yes, thank you.

OpenStudy (anonymous):

ok i can check if you want